Find the inverse of 35 modulo 11 by using extended Euclidean Algorithm


1
Expert's answer
2021-11-15T16:41:09-0500

Let us find the inverse of 35 modulo 11 by using extended Euclidean Algorithm.

Taking into account that

"35=11\cdot 3+2\\" and "11=2\cdot5+1,"

we conclude that

"1=11-2\cdot 5=11-(35-11\cdot 3)5=11\cdot 16+35(-5)."

It follows that

"11\cdot 16+35(-5)=1\ mod\ 11,"

and hence

"35(-5)=1\ mod\ 11."

Since "-5\ mod\ 11=6\ mod\ 11," we conclude that the inverse of 35 modulo 11 is 6.